OpenStudy (anonymous):

what is the integral of 3x^3/x^3+5 using integration by substitution.i need the steps

OpenStudy (anonymous):

you need to divide first

OpenStudy (anonymous):

use the terms u=x^3 and so on

OpenStudy (anonymous):

not unless it is \(3x^2\) on top. you have \(3x^3\)

OpenStudy (anonymous):

i hope that is a typo because otherwise this integral is a big fat mess

OpenStudy (anonymous):

am sorry its a typo

OpenStudy (anonymous):

it should be 3x^2

OpenStudy (anonymous):

whew then it is easy \[u=x^3+5, du =3x^2dx\] \[\int \frac{du}{u}=\ln(u)=\ln(x^3+5)\]

OpenStudy (anonymous):

thank you..i could get the answer by recognition nt by substitution thats why i asked